Find the equation of the line with y-intercept (0,8) that is perpendicular to y=1/2x=4
Answer by lwsshak3(11628)

Find the equation of the line with y-intercept (0,8) that is perpendicular to
y=1/2x-4
Standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
slope of given equation=1/2
slope of line perpendicular to given line=-2 (negative reciprocal)
Equation: y=-2x+8